1.       Diminishing Marginal Returns to labor: (2 points)

Fill in the following table and determine where diminishing marginal returns set in. Why is it that there are diminishing marginal returns to labor? Explain.

Number of Workers

Total Output

Marginal Product

0

0

1

4

2

10

3

18

4

28

5

36

6

42

7

44

8

43
Answer #1
Number of workers Total output Marginal product
0 0 -
1 4 (4-0) = 4
2 10 (10-4)=6
3 18 (18-10)= 8
4 28 (28-18) =10
5 36 (36-28)=8
6 42 (42-36)=6
7 44 (44-42)=2
8 43 (43-44)=-1

Diminishing marginal returns means when the marginal product of production decrease at a point when there is an increase in factor of production.

In this case diminishing marginal returns start to set at 5 workers with an total output of 36. From here the marginal product is decreasing which shows that there is a diminishing marginal return.

So it is clear than after increasing workers more than 4 the total oupout is increasing but the marginal product is falling.

Here diminishing marginal returns of labour occurred due to fall in marginal product of labour.